Session: Jun 2011 (session 2)
Wavelength: 6 cm


To give feedback on this experiment, click here

Experiment Feedback

Yebes (Paco Colomer at Thu Jun 02 09:20:36 2011) Minor failures: TSYS
Wrong tsys measurements on time range 4:19 - 6:29 UT.

Urumqi (Wenjun Yang at Thu Jun 02 09:35:36 2011) Minor failures: RECO
from UT 22:34 to 22:39, problems with 1pps, and from UT 02:45 to 03:20, problems with Mk5B, more details can see log file please.

Westerbork (Geert Kuper at Fri Jun 03 07:07:20 2011) Minor failures: FS
At UTD 152 00:23 UT our FieldSystem time jumps to 2010.020.21:57 . The current diskpack (TR00007) keeps on recording without stop until it is full at estimated time 01:10 UT. The problem is fixed at 02:18 UT. So no recording from 01:10 - 02:18 UT.

Medicina (Andrea Orlati at Fri Jun 03 09:22:43 2011) Failure: RECO
Mark5 fails to record at 1Gbit/sec

Torun (Kaz Borkowski at Fri Jun 03 10:28:42 2011) Minor failures:RECE HIGH
Receivers continue to be very unstable. RCP channels are practically unusable. Tsys in LCP was high in the beginning but seems to have returned to its normal temperature sometime after midnight.

Onsala (Michael Lindqvist at Mon Jun 06 05:21:45 2011) Success:
Tsys IF2 high. Otherwise no known problems.

Effelsberg (Uwe Bach at Mon Jun 06 09:49:23 2011) Success: